OpenStudy (anonymous):

Find the equation of a circle with a center at (–7, –1) where a point on the circle is (–4, 3). (x + 7)2 + (y + 1)2 = 25 (x - 7)2 + (y - 1)2 = 25 (x - 7)2 + (y - 1)2 = 5

OpenStudy (anonymous):

i think that would be the first one

OpenStudy (ranga):

The equation of a circle with center (h,k) and radius r is: (x-h)^2 + (y-k)^2 = r^2 Here (h,k) is (-7,-1) and you can find the radius by finding the distance between (-7,-1) and (-4,3).

OpenStudy (anonymous):

@ranga the radius is 5. so what do you do next?

OpenStudy (ranga):

Put h = -7, k = -1, r = 5 in (x-h)^2 + (y-k)^2 = r^2

OpenStudy (anonymous):

oh. nevermind you just plug them in. So would it be the last one?

OpenStudy (ranga):

Try again.

OpenStudy (anonymous):

the middle one?

OpenStudy (ranga):

(x - h) = (x - (-7)) = (x + 7) (y - k) = (y - (-1)) = (y + 1)

OpenStudy (anonymous):

im confused.

OpenStudy (ranga):

Put h = -7, k = -1, r = 5 in (x-h)^2 + (y-k)^2 = r^2 (x-(-7))^2 + (y-(-1))^2 = 5^2 (x+7)^2 + (y+1)^2 = 25

OpenStudy (anonymous):

okay.